OpenStudy (anonymous):

How much must you invest at one time in order to accumulate $20,000 in 15 years at 5% annual interest? S = A/[(1+ r)n] A. $5,023.07 B. $9,620.34 C. $10,023.07 D. $15,075.22

OpenStudy (ranga):

Just plug the numbers into the formula and calculate S A = 20000, r = 0.05, n = 15. Put it in S = A/[(1+ r)^n]

OpenStudy (anonymous):

i cant get it can u please show me the process ?

OpenStudy (ranga):

S = 20000 / (1 + 0.05)\[\Large S = \frac{ 20000 }{ (1 + 0.05)^{15} } = \frac{ 20000 }{ 1.05^{15} } = 9620.34\]

OpenStudy (anonymous):

in the second process you multiple 1.05 times 15 and then you divide it be 20000 right

OpenStudy (ranga):

Ignore the first line in my reply above. I forgot to delete it. 1.05 is not multiplied by 15. 1.05 is raised to 15. 15 is the exponent. \[\Large 1.05^{15} = 2.0789\] \[\Large \frac{ 20000 }{ 2.0789 } = 9620.34\]

OpenStudy (anonymous):

thank you ranga you have been really helpful:)

OpenStudy (ranga):

you are welcome.